[a4paper]article ifthen [utf8]inputenc [T1]fontenc [english]babel amsmath amssymb,amsfonts,textcomp color [top=2cm,bottom=2cm,left=2cm,right=2cm,nohead,nofoot]geometry

array supertabular hhline hyperref pdftex, colorlinks=true, linkcolor=blue, citecolor=blue, filecolor=blue, urlcolor=blue, pdftitle= [pdftex]graphicx

[gray].7 []inputenc [1]B M K R E X J o Z H N O T A *[1] 0¡¡/S/D¿¿ english Δ Δ Δ Δ Δ Δ Δ

Pin Port UART, CAN, I2C alternate
1  -  GND
2  -  +3V
3 RF2  U1RX
4 RF3  U1TX
5 RD14  -U1CTS  CN20, IC7
6 RD15  -U1RTS  CN21, IC8
7 RG2  SCL1
8 RG3  SDA1
9 RG0  CRX2
10 RG1  C2TX
11 RF4  U2RX  CN17
12 RF5  U2TX  CN18
13 RB8  -U2CTS  AN8
14 RB14  -U2RTS  AN14
15 -  GND
16 -  +5V

Δ Δ Δ [width=19.046cm,height=13.668cm]escherManual-img/escherManual-img17.pdf Δ